# PACMANN: Point Adaptive Collocation Method for Artificial Neural Networks

### v1

@@ 5      Train the PINN for $P$ iterations;
7      Determine $r^{2}(\boldsymbol{x},t)=\left(\boldsymbol{u_{t}}(\boldsymbol{x}_{r}^{i},t_{r}^{i},\boldsymbol{\theta})+\mathcal{N}[\boldsymbol{u}](\boldsymbol{x}_{r}^{i},t_{r}^{i},\boldsymbol{\theta})\right)^{2}$, the squared PDE residual over $\boldsymbol{X}$;
@@
9      Find $\nabla_{\boldsymbol{x}}r^{2}(\boldsymbol{x},t)$ and $\frac{\partial}{\partial t}r^{2}(\boldsymbol{x},t)$, the gradients of the squared PDE residual with respect to the input coordinates of the points in $\boldsymbol{X}$;
@@
11      Iteratively move the points in $\boldsymbol{X}$ according to the chosen optimization algorithm with stepsize $s$ and number of steps $T$;
@@
13      Replace points in $\boldsymbol{X}$ outside the domain with points sampled according to a uniform probability distribution;
@@
Algorithm 1 PACMANN with $P$, $s$ and $T$
Figure 1: A schematic of PACMANN with five steps of gradient ascent on a contour plot of the squared residual.
In equation (2), points are moved directly based on the gradient vector. This approach is essentially equal to applying gradient ascent. However, other gradient-based optimization algorithms can be applied. The optimization algorithms considered for PACMANN in Section 3 are listed as follows, with examples for an arbitrary variable $x$ and a function $f(x)$ to be maximized.
@@
    Gradient ascent: The collocation points are directly moved in the direction of steepest ascent. This algorithm updates variables using the principle:
    $x_{i+1}=x_{i}+s\cdot\frac{d}{dx}f(x_{i}).$
@@
    Nonlinear gradient ascent: In this algorithm, we apply a nonlinear function to the gradient ascent algorithm to scale down large gradients, preventing points from taking large steps directly out of the domain. We refer to the algorithm as nonlinear gradient ascent. In this work, we use the hyperbolic tangent function, as shown in the example:
    $x_{i+1}=x_{i}+s\cdot\tanh{\left(\frac{d}{dx}f(x_{i})\right)}.$
@@
    RMSprop [28]: RMSprop adapts the stepsize by dividing the gradient by a weighted average of previous gradients; this serves to stabilize convergence. The algorithm consists of two steps. First, a parameter $S$ is updated. This parameter consists of a weighted average of previous gradients:
    $S_{i+1}=\beta\cdot S_{i}+(1-\beta)\cdot\left(\frac{d}{dx}f(x_{i})\right)^{2}.$
    Here, $\beta$ influences the weight of the current gradient, set to $0.999$. Furthermore, $S_{0}$ is set to 00. Next, the variable $x_{i}$ is updated using:
    $x_{i+1}=x_{i}+s\cdot\frac{\frac{d}{dx}f(x_{i})}{\sqrt{S_{i+1}+\epsilon}}.$
    To prevent large steps due to small values of $S$, a small value ($10^{-8}$) is added, represented by $\epsilon$.
@@
    Momentum [29]: The momentum optimizer considers a weighted average of previous gradients at each iteration to prevent converging to local minima. First, the weighted average is computed, represented by $V$:
    $V_{i+1}=\beta\cdot V_{i}+(1-\beta)\cdot\left(\frac{d}{dx}f(x_{i})\right).$
    Here, $V_{0}$ is set to 00 and $\beta$ to $0.9$. After updating $V$, the variable $x$ is updated:
    $x_{i+1}=x_{i}+s\cdot V_{i+1}.$
@@
    Adam [26]: The Adam optimizer combines the concepts behind momentum and RMSprop. First, the value of $V$ is updated:
    $V_{i+1}=\beta_{1}\cdot V_{i}+(1-\beta_{1})\cdot\frac{d}{dx}f(x_{i}).$
    Here, $\beta_{1}$ is set to $0.9$ and $V_{0}$ to 00. Next, the parameter $S$ is updated:
    $S_{i+1}=\beta_{2}\cdot S_{i}+(1-\beta_{2})\cdot\left(\frac{d}{dx}f(x_{i})\right)^{2}.$
    Here, $\beta_{2}$ is set to $0.999$. Afterwards, an initialization bias correction is applied to $V_{i+1}$ and $S_{i+1}$:
@@
    $x_{i+1}=x_{i}+s\cdot\frac{\hat{V}_{i+1}}{\sqrt{\hat{S}_{i+1}+\epsilon}}$
    Here, $\epsilon$ is set to $10^{-8}$, preventing large steps when $\hat{S}_{i+1}$ is small.
@@
    Golden section search [30]: Golden section search is a line search method that narrows down the search interval each iteration. By searching along the direction of steepest ascent, the multidimensional optimization problem is reduced to a one-dimensional problem. In this direction, the algorithm searches in an interval $[a_{i},b_{i}]$. In the initial interval, $a_{0}$ is equal to the value $x_{0}$, for example, the x-coordinate of a collocation point. We determine $b_{0}$ using the stepsize and the gradient:
    $b_{0}=a_{0}+s\cdot\frac{d}{dx}f(x_{0}).$
    Next, $f(x)$ is evaluated at two points, $x_{l,i}$ and $x_{r,i}$, determined using:
    $x_{l,i}=a_{i}+\alpha\cdot(b_{i}-a_{i}),\quad x_{r,i}=a_{i}+\beta\cdot(b_{i}-a_{i}).$
    The name ”golden section search” refers to the golden ratio, defined as $\phi=\frac{1+\sqrt{5}}{2}$, which is incorporated in the values of $\alpha$ and $\beta$:
@@
    Else, if $f(x_{l,i})<f(x_{r,i})$, then the interval is shortened and shifted to the right:
@@
    After updating the interval to $[a_{i+1},b_{i+1}]$, the algorithm is repeated again. After the final iteration, the value of the variable $x$ is found by taking the middle point of the interval $[a_{N},b_{N}]$:
@@
    The golden ratio ensures that either $x_{l,i+1}$ ends up on $x_{r,i}$ or $x_{r,i+1}$ on $x_{l,i}$, depending on the direction of the interval shift. Since $f(x)$ has already been determined for $x_{l,i}$ and $x_{r,i}$ in the previous iteration, $f(x)$ does not have to be evaluated again for these variable values. As a result, for each iteration, the value of $f(x)$ only has to be evaluated once, lowering the computational cost.
@@ 3 Results
In this section, PACMANN is compared to collocation point sampling methods proposed in related work [17, 19]. In the following tables and figures, PACMANN is denoted by an asterisk (*) followed by the optimization method used, for example, “*Adam”. In addition, the hyperparameters of PACMANN are varied to showcase its behavior. To gather data on the prediction accuracy and computational cost, each method and set of hyperparameters is run ten times. The prediction accuracy is compared based on the mean and standard deviation of the test error, measured using the $L_{2}$ relative error. The $L_{2}$ relative error is determined as follows:
$L_{2}=\frac{\|\boldsymbol{u}_{true}-\boldsymbol{u}_{pred}\|_{2}}{\|\boldsymbol{u}_{true}\|_{2}},$ (3)
where $\boldsymbol{u}_{true}$ is the true solution and $\boldsymbol{u}_{pred}$ is the predicted solution. In the following numerical experiments, $\boldsymbol{u}_{true}$ and $\boldsymbol{u}_{pred}$ are compared over an equispaced uniform grid of 10,000 collocation points. The mean runtime of training over the ten runs serves as an indication of the computational cost of a particular sampling method. All numerical experiments were carried out with the PINNs library DeepXDE [17] using PyTorch [31] version 1.12.1 as the backend. The PINNs were trained using NVIDIA Tesla V100S GPUs on TU Delft’s high-performance computer DelftBlue [32].
For all experiments, the neural network is trained in five identical blocks of 10,000 iterations, consisting of 7000 iterations of Adam with a learning rate of $10^{-3}$ followed by 3000 iterations of L-BFGS, totaling 50,000 iterations. Throughout training, a resampling period $P$ of 50 iterations is maintained. Furthermore, resampling only occurs when Adam is set as the optimizer of the neural network. Resampling while L-BFGS is optimizer would disrupt the convergence of the algorithm due to the change in loss landscape by evaluating the PDE loss term at different collocation points. In this work, the hyperbolic tangent is used as activation function.
@@
Infrequently, the random neural network weight initialization prevents the PINN from learning the solution, which results in a test error several orders of magnitude larger than the test error obtained with other weight initializations. This has been observed for all sampling methods considered in this study and is characterized by volatile loss behavior or large static loss terms. When this occurs, the results are not considered and the PINN is trained for ten different network initializations to gather new results.
@@
We first consider the 1D Burgers’ equation:
[…]
Here, $\nu$ is the diffusion coefficient or kinematic viscosity, set at $\nu=0.01/\pi$. For this problem, the number of collocation points is set to 2500, the number of points on the boundary to 80, and the number of points on the initial condition to 160. The neural network architecture used for this example consists of four hidden layers of 64 neurons.
Sampling method $L_{2}$ relative error Mean runtime [s]
@@
*Adam 0.07% 0.05% 461
Table 1: Overview of the mean and standard deviation of the test error and the mean runtime for each sampling method for the Burgers’ equation example.
The mean and standard deviation of the test error and the mean runtime for each sampling method are given in Table 1. Out of the non-adaptive sampling methods tested, only the static uniform grid is unable to capture the solution, resulting in a high test error of $25.9\%$. The other non-adaptive methods, the static Hammersley grid, and random resampling, attain a significantly lower test error. Overall, our method, in combination with the Adam optimizer, achieves the lowest test error. It achieves a lower error than the next-best sampling method, RAR, at a slightly higher computational cost.
Table 2 compares the performance of the various optimization algorithms for our method in terms of accuracy and efficiency. We note that the performance of nonlinear gradient ascent and the RMSprop optimizers with our proposed method both achieve a competitive test error and computational cost compared to RAR.
Sampling method $L_{2}$ relative error Mean runtime [s] Hyperparameters
@@
*Gradient ascent 0.30% 0.17% 436 $10^{-6}$ 1
*Nonlinear gradient ascent 0.10% 0.06% 453 $10^{-4}$ 5
*RMSprop 0.10% 0.03% 442 $10^{-6}$ 10
*Momentum 0.18% 0.24% 448 $10^{-6}$ 5
*Adam 0.07% 0.05% 461 $10^{-5}$ 15
*Golden section search 0.34% 0.17% 460 $10^{-7}$ 5
Table 2: Overview of the mean and standard deviation of the test error and the mean runtime achieved by PACMANN for each optimization algorithm listed in Section 2, including the optimal values for the stepsize and the number of steps, for the Burgers’ equation example.
Next, we test the behavior of the different sampling methods by varying the number of collocation points from 500 to 20,000; see Figure 2(a). We observe that RAR initially reduces the test error the fastest but plateaus at a higher error than the other methods under consideration. PACMANN in combination with the Adam optimizer significantly improves the prediction accuracy as the number of collocation points is increased from 1500 to 2000, after which it slowly increases the prediction accuracy. Only at a large number of collocation points (10,000) are the adaptive sampling methods slightly outperformed by the static Hammersley grid.
@@
In the following, we compare the behavior of the optimizers listed in Section 2 when varying the hyperparameters stepsize $s$ and number of steps $T$. First, in Figure 3(a), we test the accuracy of the optimizers for different stepsizes $s$ ranging from 1 to $10^{-9}$. This figure demonstrates that the stepsize has a significant influence on the prediction accuracy achieved by PACMANN. We note that, depending on the optimizer used, a different stepsize is optimal, such as $10^{-6}$ for momentum or $10^{-5}$ for Adam. Furthermore, the behavior of the optimizers at stepsizes near $1$ is split into two groups, with Adam and RMSprop gaining accuracy as the stepsize is increased, while others continue to lose accuracy. This phenomenon is explained by the number of collocation points that leave the domain while points are moved by PACMANN. At these stepsizes all 2500 collocation points exit the domain when using Adam or RMSprop. Since PACMANN uses a uniform probability distribution to determine the location of the replacement collocation points, the test error approaches the accuracy of random resampling ($0.40\%\pm 0.35\%$). In contrast, when applying the other optimization algorithms with these stepsizes, only a small portion of the collocation points exit the domain, about ten on average. These few points are not significant enough for the random resampling of points to affect the test error achieved. The difference in the number of points that exit the domain may be explained by the relatively large optimal stepsize and number of steps for RMSprop and Adam compared to the other optimization algorithms.
@@
Figure 4: Mean of the test error and the training time for PACMANN with Adam for a varying (a) number of collocation points and (b) number of steps for the Burgers’ equation example.
Furthermore, Figures 4(a) and 4(b) depict the prediction accuracy and computational cost of PACMANN with Adam for varying numbers of collocation points and steps. Importantly, we point out that the accuracy of Adam with five steps is nearly the same as its accuracy at 15 steps, however, repeating our method five times is less expensive. Thus, we recommend considering taking fewer steps to save on computational cost in more complex problems.
Figure 5: Location of the 2500 collocation points (a) before training and after training with (b) RAR, (c) RAD, and (d) PACMANN with Adam for the Burgers’ equation example.
Next, we compare the distribution of collocation points after training. Figure 5(a) shows the locations before training when the collocation points are laid out based on the Hammersley sequence. Figures 5(b), 5(c) and 5(d) show the locations of the collocation points after training for RAR, RAD, and PACMANN with Adam, respectively. While RAR creates one large cluster of points at the steepest region of the solution, RAD and our method tend to create several smaller clusters. In contrast to RAD, our method also forms clusters of points in regions with typically lower residuals, indicative of local maxima of the squared residual. We note the similarity between the shape of the clusters and the solution.
@@
In the following example, we consider the 1D Allen-Cahn equation:
[…]
Let the diffusion coefficient $d=0.001$. Similar to the previous Burgers’ equation problem, the number of collocation points is set to 2500, the number of boundary points to 80, and the initial condition points to 160. The network architecture used for this example consists of four hidden layers of 64 neurons.
Sampling method $L_{2}$ relative error Mean runtime [s]

### v2

@@ 5  Train the PINN for $P$ iterations;
7  Determine $r^{2}(\boldsymbol{x},t)=\left(\boldsymbol{u_{t}}(\boldsymbol{x}_{r}^{i},t_{r}^{i},\boldsymbol{\theta})+\mathcal{N}[\boldsymbol{u}](\boldsymbol{x}_{r}^{i},t_{r}^{i},\boldsymbol{\theta})\right)^{2}$, the squared PDE residual over $\boldsymbol{X_{r}}$;
@@
9  Find $\nabla_{\boldsymbol{x}}r^{2}(\boldsymbol{x},t)$ and $\frac{\partial}{\partial t}r^{2}(\boldsymbol{x},t)$, the gradients of the squared PDE residual with respect to the input coordinates of the points in $\boldsymbol{X}_{r}$;
@@
11  Iteratively move the points in $\boldsymbol{X}_{r}$ according to the chosen optimization algorithm with stepsize $s$ and number of steps $T$;
@@
13  Replace points in $\boldsymbol{X}_{r}$ outside the domain with points sampled according to a uniform probability distribution;
@@
Algorithm 1 PACMANN with a given optimization algorithm, $P$, $s$, and $T$
Figure 1: A schematic of PACMANN with four steps of gradient ascent on a contour plot of the squared residual.
In this work, we propose the Point Adaptive Collocation Method for Artificial Neural Networks, which uses the gradient of the squared residual as guiding information to gradually move collocation points toward areas of high residuals; see Algorithm 1 and Figure 1. In particular, instead of the minimization problem in Equation 2, we consider the following min-max problem for training the model parameters $\boldsymbol{\theta}$:
$\boldsymbol{\theta}^{*}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\left[\lambda_{ic}\mathcal{L}_{ic}\left(\boldsymbol{\theta}\right)+\lambda_{bc}\mathcal{L}_{bc}\left(\boldsymbol{\theta}\right)+\lambda_{ref}\mathcal{L}_{ref}\left(\boldsymbol{\theta}\right)+\lambda_{r}\max_{\boldsymbol{X_{r}}\subset\mathcal{D}}\mathcal{L}_{r}\left(\boldsymbol{X_{r}},\boldsymbol{\theta}\right)\right].$ (7)
Here, only the collocation points $X_{r}=\{(\boldsymbol{x}_{r}^{i},t_{r}^{i})\}_{i=1}^{N_{r}}$ are moved, while points such as those sampled along the boundary are fixed in place throughout training. This approach ensures that a sufficient number of points are placed along the initial and boundary conditions to accurately compute the respective loss terms.
Remark.
The min-max formulation for the residual loss $\mathcal{L}_{r}$ in Equation 7 can also be interpreted as optimizing the supremum norm in $\mathcal{D}$. In our numerical experiments, the neural network function will be smooth, due to the use of the hyperbolic activation function, such that the maximum exists in $\mathcal{D}$. Then, we obviously have that
$\left\|\hat{\boldsymbol{u_{t}}}+\mathcal{N}[\hat{\boldsymbol{u}}]\right\|_{L^{\infty}(\mathcal{D})}^{2}$ $=\max_{\left(\boldsymbol{x}_{r},t_{r}\right)\in\mathcal{D}}\left(\hat{\boldsymbol{u_{t}}}\left(\boldsymbol{x}_{r},t_{r}\right)+\mathcal{N}[\hat{\boldsymbol{u}}]\left(\boldsymbol{x}_{r},t_{r}\right)\right)^{2}=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\max_{\left(\boldsymbol{x}_{r},t_{r}\right)\in\mathcal{D}}\left(\hat{\boldsymbol{u_{t}}}\left(\boldsymbol{x}_{r},t_{r}\right)+\mathcal{N}[\hat{\boldsymbol{u}}]\left(\boldsymbol{x}_{r},t_{r}\right)\right)^{2}$
$=\max_{\boldsymbol{X}_{r}\subset\mathcal{D}}\left[\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\hat{\boldsymbol{u_{t}}}\left(\boldsymbol{x}_{r}^{i},t_{r}^{i}\right)+\mathcal{N}[\hat{\boldsymbol{u}}]\left(\boldsymbol{x}_{r}^{i},t_{r}^{i}\right)\right)^{2}\right]=\max_{\boldsymbol{X}_{r}\subset\mathcal{D}}\mathcal{L}_{r}\left(\boldsymbol{X_{r}},\boldsymbol{\theta}\right),$
where we have omitted the dependence on the neural network parameters $\theta$ for the sake of brevity. In practice, since we never actually attain the maximum, the loss will be significantly lower than the supremum norm, due to the averaging across $\mathcal{D}$.
In the PACMANN algorithm, $N_{r}$ collocation points are first sampled using a uniform sampling method, such as an equispaced uniform grid or the Hammersley sequence [29, pp. 31–36]. The PINN is then trained on this set of collocation points for a number of $P$ iterations. The number $P$ is a hyperparameter of the method determining the resampling period, that is, the period after which the collocation points are resampled. Next, the training iteration is paused and the gradients of the squared residual $r^{2}(\boldsymbol{x},t)$ with respect to the input coordinates, given by $\nabla_{\boldsymbol{x}}r^{2}(\boldsymbol{x},t)$ and $\frac{\partial}{\partial t}r^{2}(\boldsymbol{x},t)$, are determined for each collocation point $\left(\boldsymbol{x},t\right)\in\boldsymbol{X}_{r}$. The collocation points are subsequently moved in the direction of increasing residual based on
[…]
Here, $s$ is a hyperparameter that determines the stepsize of each move; in the machine learning community, this parameter is also often called a learning rate. Since the neural network parameters $\theta$ are kept constant during the iteration in Equation 8, the residual landscape is static and the residual gradient can be determined again for the new location of the collocation point, allowing the process to be repeated. Therefore, collocation points can be moved several times, given by the hyperparameter number of steps $T$. If a point moves outside the domain, it is removed from the set $\boldsymbol{X}_{r}$ and a replacement point is sampled in the domain based on a uniform probability distribution. Next, the neural network is trained for another $P$ iterations, after which the process of moving collocation points for $T$ iterations is repeated.
In Equation 8, points are moved directly based on the gradient vector. This approach is essentially equal to applying gradient ascent. However, other gradient-based optimization algorithms can be applied. The optimization algorithms considered for PACMANN in Section 3 are listed as follows. We also provide the iteration rule for an arbitrary variable $x$ and a function $f(x)$ to be maximized. Other optimization algorithms are also applicable, but we focus on the following algorithms in our numerical experiments in Section 3 and A.
@@
    Gradient ascent: The collocation points are directly moved in the direction of steepest ascent, that is, in the direction of the gradient. This algorithm updates variables using the formula:
    $x_{i+1}=x_{i}+sf^{\prime}(x_{i}).$
@@
    Nonlinear gradient ascent: In this algorithm, we apply a nonlinear function to the gradient ascent algorithm to scale down large gradients, preventing points from taking large steps directly out of the domain. We refer to the algorithm as nonlinear gradient ascent. In this work, we use the hyperbolic tangent function, as follows:
    $x_{i+1}=x_{i}+s\tanh{\left(f^{\prime}(x_{i})\right)}.$
@@
    RMSprop: Root Mean Square Propagation (RMSprop) [30] adapts the stepsize by dividing the gradient by a weighted average of previous gradients; this serves to stabilize convergence. The algorithm consists of two steps. First, a parameter $S$ is updated. This parameter consists of a weighted average of previous gradients:
    $S_{i+1}=\beta S_{i}+(1-\beta)\left(f^{\prime}(x_{i})\right)^{2}.$
    Next, the variable $x$ is updated using:
    $x_{i+1}=x_{i}+s\frac{f^{\prime}(x_{i})}{\sqrt{S_{i+1}+\epsilon}}.$
    To prevent large steps due to small values of $S$, a small value is added, represented by $\epsilon$.
@@
    Momentum: The momentum optimizer [31] considers a weighted average of previous gradients at each iteration to prevent converging to local minima. First, the weighted average, $V$, is computed:
    $V_{i+1}=\beta V_{i}+(1-\beta)\left(f^{\prime}(x_{i})\right).$
    After updating $V$, the variable $x$ is updated:
    $x_{i+1}=x_{i}+sV_{i+1}.$
@@
    Adam: The Adaptive moments (Adam) [28] optimizer combines the concepts behind RMSprop and momentum. First, the value of $V$ is updated:
    $V_{i+1}=\beta_{1}V_{i}+(1-\beta_{1})f^{\prime}(x_{i}).$
    Next, the parameter $S$ is updated:
    $S_{i+1}=\beta_{2}S_{i}+(1-\beta_{2})\left(f^{\prime}(x_{i})\right)^{2}.$
    Afterwards, an initialization bias correction is applied to $V_{i+1}$ and $S_{i+1}$:
@@
    $x_{i+1}=x_{i}+s\frac{\hat{V}_{i+1}}{\sqrt{\hat{S}_{i+1}+\epsilon}}$
    Here, a small regularization parameter $\epsilon$ is included to prevent large steps when $\hat{S}_{i+1}$ is small.
@@
    Golden section search: Golden section search [32, pp. 39–42] is a line search method that narrows down the search interval each iteration. By searching along the direction of steepest ascent, the multidimensional optimization problem is reduced to a one-dimensional problem. In this direction, the algorithm searches in an interval $a_{i},b_{i}]$. In the initial interval, $a_{0}$ is equal to the value $x_{0}$, for example. We determine $b_{0}$ using the stepsize and the gradient:
    $b_{0}=a_{0}+sf^{\prime}(x_{0}).$
    Next, $f(x)$ is evaluated at two points, $x_{l,i}$ and $x_{r,i}$, determined using
    $x_{l,i}=a_{i}+\alpha(b_{i}-a_{i}),\quad x_{r,i}=a_{i}+\beta(b_{i}-a_{i}).$
    The name “golden section search” refers to the golden ratio, defined as $\phi=\frac{1+\sqrt{5}}{2}$, which is incorporated in the values of $\alpha$ and $\beta$:
@@
    Otherwise, if $f(x_{l,i})<f(x_{r,i})$, then the interval is shortened and shifted to the right:
@@
    After updating the interval to $a_{i+1},b_{i+1}]$, the algorithm is repeated again. After the final iteration, the value of the variable $x$ is found by taking the middle point of the interval $a_{N},b_{N}]$:
@@
    The golden ratio ensures that either $x_{l,i+1}$ ends up on $x_{r,i}$ or $x_{r,i+1}$ on $x_{l,i}$, depending on the direction of the interval shift. Since $f(x)$ has already been determined for $x_{l,i}$ and $x_{r,i}$ in the previous iteration, $f(x)$ does not have to be evaluated again for these variable values. As a result, for each iteration, the value of $f(x)$ only has to be evaluated once, which is beneficial in terms of the computational cost.
@@ 3 Results
In this section, we evaluate the performance of PACMANN in terms of accuracy and computational cost across various PDE examples, including simple model problems, an inverse problem, a problem defined on an irregular-shaped domain, and two high-dimensional problems. Furthermore, we vary the hyperparameters of PACMANN to showcase its behavior and compare our method with other collocation point sampling methods proposed in the related works [18, 20].
To gather data on the prediction accuracy and computational cost, each method and set of hyperparameters is run ten times with varying random seeds. The prediction accuracy is compared based on the mean and standard deviation across the ten runs of the test error, measured using the $L_{2}$ relative error. The $L_{2}$ relative error is determined as follows:
$\varepsilon_{L_{2}}\coloneqq\frac{\|\boldsymbol{u}_{ref}-\boldsymbol{u}_{pred}\|_{2}}{\|\boldsymbol{u}_{ref}\|_{2}},$ (9)
where $\boldsymbol{u}_{ref}$ is the reference solution, which is either an analytical or numerical solution depending on the problem, and $\boldsymbol{u}_{pred}$ is the predicted solution. When an analytical solution is available, we additionally consider the prediction accuracy using the $H^{1}$ semi-norm
$\varepsilon_{H^{1}}\coloneqq\frac{|\boldsymbol{u}_{ref}-\boldsymbol{u}_{pred}|_{1}}{|\boldsymbol{u}_{ref}|_{1}},\quad\text{where}\quad|v|_{1}\coloneqq\|\nabla v\|_{2},$ (10)
to measure errors in the first derivatives of the solution. To compare $\boldsymbol{u}_{ref}$ and $\boldsymbol{u}_{pred}$ we employ an equispaced uniform grid of 10 000 collocation points. The mean runtime of training over the ten runs serves as an indication of the computational cost of a particular sampling method. Our code is based on the PINNs library DeepXDE [18] using PyTorch [33] version 1.12.1 as the backend. It is publicly available on GitHub at https://github.com/CoenVisser/PACMANN. The models were trained using NVIDIA Tesla V100S GPUs on TU Delft’s high-performance computer DelftBlue [34].
For all experiments, the training is split into five phases of 10 000 iterations, consisting of 7000 iterations of Adam with a learning rate of $10^{-3}$ followed by 3000 iterations of the Limited-memory Broyden-Fletcher-Goldfarb-Shanno (L-BFGS) algorithm [35], totaling 50 000 iterations. Throughout training, a resampling period $P$ of 50 iterations is maintained. Furthermore, we only resample the collocation points when training the neural network parameters with Adam. Resampling while training with the L-BFGS optimizer would disrupt the convergence of the algorithm due to the change in loss landscape by evaluating the PDE loss term at different collocation points. In this work, in all experiments, the hyperbolic tangent is used as the activation function. In addition, the hyperparameter settings used for RMSprop, momentum, and Adam with PACMANN are given in Table 1. Note that we have not conducted an extensive study varying the neural network architecture. Instead, we have used the architectures provided in the corresponding test cases of DeepXDE, assuming that these were already optimized appropriately.
Optimizer $\boldsymbol{\beta}$ $\boldsymbol{V_{0}}$ $\boldsymbol{S_{0}}$ $\boldsymbol{\epsilon}$
RMSprop 0.999 - 0 $10^{-8}$
Momentum 0.9 0 - $10^{-8}$
Adam $\beta_{1}=0.9$, $\beta_{2}=0.999$ 0 0 $10^{-8}$
Table 1: Hyperparameter settings for the RMSprop, momentum, and Adam optimizers.
@@
Infrequently, the random neural network weight initialization prevents the PINN from learning the solution, which results in a test error several orders of magnitude larger than the test error obtained with other weight initializations. This has been observed for all sampling methods considered in this study and is characterized by volatile loss behavior or large static loss terms. When this occurs, the corresponding training run is repeated with a different random seed.
@@
Sampling method $L_{2}$ relative error
Mean
runtime [s]
@@
PACMANN-Adam 0.07% 0.05% 461
Table 2: Overview of the mean and standard deviation of the test error and the mean runtime for each sampling method for the Burgers’ equation. The best result in each column is marked in boldface.
PACMANN optimizer $L_{2}$ relative error
Mean
runtime [s]
Hyperparameters
@@
Gradient ascent 0.30% 0.17% 436 $10^{-6}$ 1
Nonlinear gradient ascent 0.10% 0.06% 453 $10^{-4}$ 5
RMSprop 0.10% 0.03% 442 $10^{-6}$ 10
Momentum 0.18% 0.24% 448 $10^{-6}$ 5
Adam 0.07% 0.05% 461 $10^{-5}$ 15
Golden section search 0.34% 0.17% 460 $10^{-7}$ 5
Table 3: Overview of the mean and standard deviation of the test error and the mean runtime achieved by PACMANN for the Burgers’ equation with the optimization methods listed in Section 2. The best result in each column is marked in boldface. For each optimization method, we report the stepsize and number of steps that achieve the lowest test error.
@@
Figure 4: Mean of the test error and the runtime for PACMANN with Adam for a varying (a) number of collocation points and (b) number of steps for the Burgers’ equation example.
Figure 5: Location of the 2500 collocation points (a) before training, and after training with (b) RAR, (c) RAD, and (d) PACMANN with Adam for the Burgers’ equation example. The color indicates the values of the predicted solution.
Figure 6: Location of the collocation points (a) mid-training at 25 000 iterations, and (b) after training using PACMANN with Adam for the Burgers’ equation example. The color indicates the values of the squared residual.
We first consider the one-dimensional Burgers’ equation:
[…]
Here, $\nu$ is the diffusion coefficient or kinematic viscosity, set at $\nu=0.01/\pi$. For this problem, we employ 2500 collocation points, 80 boundary points, and 160 initial condition points. The neural network architecture used for this example consists of four hidden layers of 64 neurons. To compute the $L_{2}$ relative error, we compare the network prediction to a numerical solution generated using a spectral solver by Raissi et al. [6].
The mean and standard deviation of the test error and the mean runtime for each sampling method are given in Table 2. Out of the non-adaptive sampling methods tested, only the static uniform grid is unable to capture the solution satisfactorily, resulting in a high test error of $25.9\%$. The other non-adaptive methods, the static Hammersley grid and random resampling, attain a significantly lower test error. Overall, our method, in combination with the Adam optimizer and a stepsize of $10^{-5}$, achieves the lowest test error. It achieves a lower error than the next-best sampling method, RAR, at a slightly higher computational cost.
Table 3 compares the performance of the various optimization algorithms for PACMANN in terms of accuracy and efficiency. We note that the nonlinear gradient ascent and the RMSprop optimizers with our proposed method both achieve a competitive test error and computational cost compared to the second best approach in Table 2, that is, RAR.
Next, we test the behavior of the different sampling methods by varying the number of collocation points from 500 to 20 000; see Figure 2(a). We observe that RAR initially reduces the test error the fastest, but plateaus at a higher error than the other methods under consideration. PACMANN in combination with the Adam optimizer significantly improves the prediction accuracy as the number of collocation points is increased from 1500 to 2000, after which it slowly increases the prediction accuracy. Only at a large number of collocation points (10 000) are the adaptive sampling methods slightly outperformed by the static Hammersley grid.
@@
In the following, we compare the behavior of the optimizers listed in Section 2 when varying the hyperparameters stepsize $s$ and number of steps $T$. First, in Figure 3(a), we test the accuracy of the optimizers for different stepsizes $s$ ranging from $1$ to $10^{-9}$. This figure demonstrates that the stepsize has a significant influence on the prediction accuracy achieved by PACMANN. We note that, depending on the optimizer used, a different stepsize is optimal, such as $10^{-6}$ for momentum or $10^{-5}$ for Adam. Furthermore, the behavior of the optimizers at stepsizes near $1$ is split into two groups, with Adam and RMSprop gaining accuracy as the stepsize is increased, while others continue to lose accuracy. This phenomenon is explained by the number of collocation points that leave the domain while points are moved by PACMANN. At these large stepsizes, all 2500 collocation points exit the domain when using Adam or RMSprop. Since PACMANN uses a uniform probability distribution to determine the location of the replacement collocation points, the test error approaches the accuracy of random resampling ($0.40\%\pm 0.35\%$); cf. Table 2. In contrast, when applying the other optimization algorithms with these stepsizes, only a small portion of the collocation points exit the domain, of order $\mathcal{O}(10)$. These few points are not sufficient for the random resampling to significantly affect the test error achieved. The difference in the number of points that exit the domain may be explained by the relatively large optimal stepsize and number of steps for RMSprop and Adam compared to the other optimization algorithms.
@@
Furthermore, Figures 4(a) and 4(b) depict the prediction accuracy and computational cost of PACMANN with the Adam optimizer for varying numbers of collocation points and steps. Figure 4(a) demonstrates that increasing the number of collocation points reduces the $L_{2}$ relative error before reaching a plateau, beyond which the computational cost rises steeply without a further increase in accuracy. Importantly, we point out that the accuracy of Adam with five steps is nearly the same as its accuracy at 15 steps, see  Figure 4(b). Thus, we recommend taking fewer steps to save on computational cost in more complex problems.
Next, we compare visually the distribution of collocation points after training. Figure 5(a) shows the locations before training when the collocation points are laid out based on the Hammersley sequence. Figures 5(b), 5(c) and 5(d) show the locations of the collocation points after training for RAR, RAD, and PACMANN with Adam, respectively. While RAR clusters the points at the steepest region of the solution, RAD and our method tend to create several smaller clusters. In contrast to RAD, our method also forms clusters of points in regions with typically lower residuals, see  Figure 6(a), indicative of local maxima of the squared residual. Figure 6(b) shows that, after training the PINN, these previously observed local maxima have reduced. For our approach, we note the similarity between the shape of the clusters and the solution itself.
@@
Sampling method $L_{2}$ relative error
Mean
runtime [s]